Cant solve this:
_ - _ x _ / _ = 78
using 84, 54, 9, 1
Fill in the blanks
1
Expert's answer
2011-10-06T12:11:57-0400
It's possible to fill in the blanks using 84, 54, 9, 1 like this:

84-54*1/9 = 78
(84-6=78)